On January 1, Year 1 XYZ Company paid $88,000 cash to purchase a truck. The truck has a $4,000 salvage value and a 4 year useful life. XYZ uses double declining balance depreciation. How much depreciation expense would XYZ report on its Year 2 income statement?

  • $21,000

  • $22,000

  • $16,000

  • $44,000
Answer #1

Double declining rate = 200/useful life = 200/4 = 50%

Depreciation expense for year 1

= 88,000*50% = 44,000

Depreciation expense for year 2

= (88,000-44,000)*50%

= 22,000

Option B